# Multiplicity of powers of path ideals of a line graph

### v1

@@ 2. Associated primes of $I_{t}(L_{n})^{s}$
Let $G=\{a_{1},a_{2},\ldots,a_{2p}\}$. Then $n\geq a_{2p}=(t+1)p\geq n+t-1$ and $a_{2p-1}=t(p-1)+p<n-t+1$, and so it is routine to verify that $G\in C_{n,t}$. This proves $D\geq 2p$. To prove $D\leq 2p$, we assume on the contrary that there exists $F=\{i_{1},\ldots,i_{r}\}\in C_{n,t}$ with $1\leq i_{1}<\cdots<i_{r}\leq n$ such that $r>2p$. Then $i_{r-1}\geq i_{2p}\geq a_{2p}\geq n-t+1$. This is contradicted to condition (5). Hence, $D=2p$.
Case 2: If $n=(t+1)p+t$, then $n-t+1=(t+1)p+1=a_{2p+1},$ and one easily checks that $\{a_{1},\ldots,a_{2p+1}\}$ belongs to $C_{n,t}$. It follows that $D=2p+1$, by using the same argument as in case 1.
The theory of Alexander duality plays an important role in the study of squarefree monomial ideals. Let $I=I(\Delta)$ be a squarefree monomial ideal. The Alexander dual of $I$ is defined to be the monomial ideal
$I^{\vee}=(x_{C}\colon\;C\mbox{ is a vertex cover of }\Delta).$
A well-known result states that $S/I$ is Cohen-Macaulay if and only if $I^{\vee}$ has a linear resolution, (see e.g. [10, Theorem 8.1.9]). Moreover, according to [10, Theorem 8.2.20], we know that $S/I$ is sequentially Cohen-Macaulay if and only if $I^{\vee}$ is componentwise linear. Lastly, for any squarefree monomial ideal $I$ of $S$, we have $\operatorname{pd}(S/I)=\operatorname{reg}(I^{\vee})$, as stated in [10, Proposition 8.1.10]. Here, $\operatorname{pd}$ refers to the projective dimension and $\operatorname{reg}$ refers to the (Castelnuovo-Mumford) regularity. To learn more about Alexander duality, one can refer to Chapter 8 of the book [10].
The following result recovers [13, Theorem 4.1], which is also proven in a different manner in [4].
@@
In [9, Corollary 5.4], it was shown that if $\Delta$ is a simplicial forest then $S/I(\Delta)$ is sequentially Cohen-Macaulay. Applying this result and in view of Lemma 2.2, we can conclude that $S/I_{t}(L_{n})$ is sequentially Cohen-Macaulay.
By [10, Theorem 8.2.20], we know $I_{t}(L_{n})^{\vee}$ is componentwise linear. Therefore, using [10, Corollary 8.2.14], we have $\operatorname{reg}(I_{t}(L_{n})^{\vee})=\operatorname{Deg}(I_{t}(L_{n})^{\vee})$. Furthermore, according to [10, Proposition 8.1.10], we have $\operatorname{pd}(S/I_{t}(L_{n}))=\operatorname{Deg}(I_{t}(L_{n})^{\vee})$. The result now follows from Corollary 2.9.
@@ 3. Multiplicity
In this section, we present the main result of this paper, where we provide an explicit expression for the multiplicity of powers of $I_{t}(L_{n}).$ We begin with a result regarding the multiplicity of square-free monomial ideals.
@@
This follows directly from [10, Corollary 6.2.3].
@@
Assume that $n=at+b$ with $a\geq 1$ and $0\leq b<t$. Then
$\operatorname{height}(I_{t}(L_{n}))=a.$
@@
We next describe the prime ideals of $I_{t}(L_{n})$ of height $a$.
@@
Assume $n=at+b$ with $a\geq 1$ and $0\leq b<t$. Let $\operatorname{\mathfrak{p}}=(x_{i_{1}},\ldots,x_{i_{a}})$ with $1\leq i_{1}<\cdots<i_{a}\leq n$. Then $\operatorname{\mathfrak{p}}\in\mathrm{ass}(I_{t}(L_{n}))$ if and only if the following conditions are satisfied:
@@
Assume that the elements of $F=\{i_{1},\ldots,i_{a}\}$ satisfies (1), (2). We only need to prove $F\in C_{n,t}$, as defined in Notation 2.6. First, it is easy to see that $i_{1}\leq t$ and $i_{2}\geq(2-1)t+b+1\geq t+1$. Next, we observe that $i_{a-1}\leq(a-1)t=n-t-b<n-t+1$ and $i_{a}\geq(a-1)t+b+1=n-t+1$. Finally, for $j=1,\ldots,a-2$, we see that
$i_{j+2}-i_{j}\geq(j+1)t+b+1-jt=t+b+1\geq t+1.$
Thus, $F\in C_{n,t}$ and so $\operatorname{\mathfrak{p}}\in\mathrm{ass}(I_{t}(L_{n}))$ by Theorem 2.7.
Conversely, first note that $\{i_{1},\ldots,i_{a}\}$ is a minimal vertex cover of $\Delta_{n,t}$. Since the $a$ facets $F_{1},F_{t+1},\ldots,F_{(a-1)t+1}$ of $\Delta_{n,t}$ are pairwise disjoint, it follows that $i_{j}\in F_{j}$ for $j=1,\ldots,a$, and thus $(j-1)t+b+1\leq i_{j}\leq jt$. This implies (1) has to be satisfied. As to condition (2), it follows immediately from Theorem 2.7.
To determine the multiplicity of $S/I_{t}(L_{n})$, in view of Lemma 3.1, we need to count the number of the sequences $(i_{1},\ldots,i_{a})$ that satisfy condition (1) and (2), as stated in Lemma 3.3. This is a purely combinatorial question, which will be solved using the following two results.
@@
Given $a\geq 1$ and $t\geq 1$, let $T_{a,t}$ denote the set of integral sequences $i_{1}<\ldots<i_{a}$ such that
@@
$|T_{a,t}|=\begin{pmatrix}a+t-1\\ a\end{pmatrix}.$
@@
We proceed by induction on $t$. If $t=1$, the result follows easily. Assume now that $t\geq 2$. For each sequence $(i_{1},i_{2},\ldots,i_{a})\in T_{a,t}$, we define
$\tau(i_{1},\ldots,i_{a}):=\operatorname{min}\{j\in[a]\colon\;i_{j}=(j-1)t+1\}.$
By convention, we set $\tau(i_{1},\ldots,i_{a})=a+1$ if $i_{j}\neq(j-1)t+1$ for all $j=1,\ldots,a$. Thus, $\tau(i_{1},\ldots,i_{a})\in\{1,\ldots,a,a+1\}$. We make the following easy observation:
if $\tau(i_{1},\ldots,i_{a})=k<a$, then $i_{j}=(j-1)t+1$ for $j=k,\ldots,a$.
For $k\in[a+1]$, we define
$T_{a,t}^{k}=\{(i_{1},i_{2},\ldots,i_{a})\in T_{a,t}\colon\;\tau(i_{1},\ldots,i_{a})=k\}.$
One easily see that $|T_{a,t}^{1}|=1$ and $|T_{a,t}^{2}|=|\{i_{1}:2\leq i_{1}\leq t\}|=t-1$.
In general, for $k\in[a+1]$, we let $T_{k}$ denote the set of all sequences $i_{1}<i_{2}<\ldots<i_{k-1}$ such that $(j-1)t+2\leq i_{j}\leq jt$ for $j=1,\ldots,k-1,$ and $i_{j+1}-i_{j}\leq t$ for $j=1,\ldots,k-2.$ It follows from the above observation that $|T_{a,t}^{k}|=|T_{k}|$ for $k=1,\ldots,a+1$.
@@
$(i_{1},\ldots,i_{a})\longmapsto(i_{1}^{\prime},\ldots,i_{a}^{\prime})$
establishes a bijection between $T_{k}$ and $T_{t-1,k-1}$. Hence, $|T_{k}|$ is equal to the number of $T_{t-1,k-1}$. By inductive hypothesis, we have $|T_{t-1,k-1}|=\begin{pmatrix}k+t-3\\ k-1\end{pmatrix}.$ Hence, it follows that
$|T_{a,t}|$ $=\mathop{\sum}\limits_{k=1}^{a+1}|T^{k}_{a,t}|=\mathop{\sum}\limits_{k=1}^{a+1}\begin{pmatrix}k+t-3\\ k-1\end{pmatrix}$
$=\begin{pmatrix}t-2\\ 0\end{pmatrix}+\begin{pmatrix}t-1\\ 1\end{pmatrix}+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}a+t-3\\ a\end{pmatrix}+\begin{pmatrix}a+t-2\\ a\end{pmatrix}$
$=(\begin{pmatrix}t-1\\ 0\end{pmatrix}+\begin{pmatrix}t-1\\ 1\end{pmatrix})+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}a+t-3\\ a\end{pmatrix}+\begin{pmatrix}a+t-2\\ a\end{pmatrix}$
$=\begin{pmatrix}t\\ 1\end{pmatrix}+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}a+t-3\\ a\end{pmatrix}+\begin{pmatrix}a+t-2\\ a\end{pmatrix}$
$=\begin{pmatrix}t+a-2\\ a-1\end{pmatrix}+\begin{pmatrix}a+t-2\\ a\end{pmatrix}$
$=\begin{pmatrix}a+t-1\\ a\end{pmatrix}.$
@@
Given $n\geq t\geq 1$ and assume $n=at+b$ with $10\leq b\leq t-$ 1. Let $X_{n,t}$ denote the set of integral sequences $i_{1}<\ldots<i_{a}$ such that
@@
$|X_{n,t}|=\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}.$
@@
Set $i_{j}^{\prime}=i_{j}-bj$. Then, we have for $1\leq j\leq a$,
$(j-1)t+b+1\leq i_{j}\leq jt\Longleftrightarrow(j-1)(t-b)+1\leq i_{j}^{\prime}\leq j(t-b),$
and for $11\leq j\leq a-$ 1,
$i_{j+1}-i_{j}\leq t\Longleftrightarrow i_{j+1}^{\prime}-i_{j}^{\prime}\leq t-b.$
Thus, the map: $(i_{1},\ldots,i_{a})\longmapsto(i_{1}^{\prime},\ldots,i_{a}^{\prime})$ establishes a bijection between $X_{n,t}$ and $T_{a,t-b}$. This implies
$|X_{n,t}|=|T_{a,t-b}|=\begin{pmatrix}a+t-b-1\\ a\end{pmatrix},$
@@
Assume $n=at+b$ with $a\geq 1$ and $0\leq b<t$. Then
$\operatorname{mult}(S/I_{t}(L_{n}))=\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}.$
@@
This result is a combination of Lemma 3.1, Lemma 3.3, and Corollary 3.5.
Based on Proposition 3.6, our main result can be obtained through an induction process. The crucial step in the proof is to identify suitable short exact sequences. In order to do this, we require the following lemmas.
@@
$(I+J):K=(I:K)+(J:K)=(I:K)+J.$
This fact is also useful in the following proof.
@@
The equality that $B_{i}=(I_{t}(L_{n-i})^{s},x_{n-i+1})$ follows easily. Let us show that $A_{i}:x_{n-i+1}=A_{i+1}$. Since
@@
Assume $n=at+b$, with $a\geq 1$ and $0\leq b<t$. Then, for all $s\geq 1$, we have
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\begin{pmatrix}s+a-1\\ s-1\end{pmatrix}\begin{pmatrix}a+t-b-1\\ a\end{pmatrix}.$
@@
It follows that $\mathbf{Q}(S/I_{t}(L_{t})^{s},z)=\sum_{i=0}^{st-1}z^{i}$ and so $\operatorname{mult}(S/I_{t}(L_{t})^{s})=\mathbf{Q}(S/I_{t}(L_{t})^{s},1)=st.$ This proves the result holds when $n=t$. Now, assume $s>1$ and $n>t$. To begin, we write $n=at+b$, where $a\geq 1$ and $0\leq b<t$. There are two cases to consider.
Case 1: $b>0$. Consider the following three short exact sequences of graded $S$-modules:
$0\longrightarrow\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}}[-t]\longrightarrow\frac{S}{I_{t}(L_{n})^{s}}\longrightarrow\frac{S}{(I_{t}(L_{n})^{s},u_{n-t+1})}\longrightarrow 0,$
@@
Note that $(I_{t}(L_{n})^{s},u_{n-t+1})=(I_{t}(L_{n-1})^{s},u_{n-t+1}),$ it follows from the additivity of Hilbert series that
(1) $\begin{split}&\operatorname{HS}(\frac{S}{I_{t}(L_{n})^{s}},z)=z^{t}\operatorname{HS}(\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}},z)+\operatorname{HS}(\frac{S}{I_{t}(L_{n-1})^{s}},z)\\ &-z^{1+t}\operatorname{HS}(\frac{S}{I_{t}(L_{n-1})^{s}:x_{n-t}u_{n-t+1}},z)\quad-z^{t}\operatorname{HS}(\frac{S}{(I_{t}(L_{n-1})^{s}:u_{n-t+1},x_{n-t})},z).\end{split}$
Since $I_{t}(L_{n})^{s}:u_{n-t+1}=I_{t}(L_{n})^{s-1}$ by Lemma 3.7, we have $\dim\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}}=n-a$ and thus
(2) $\operatorname{HS}(\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}},z)=\frac{1}{(1-z)^{n-a}}\mathbf{Q}(\frac{S}{I_{t}(L_{n})^{s-1}},z).$
Denote $S_{j}=K[x_{1},\ldots,x_{j}]$ for $j=1,\ldots,n$. Since $n-1=at+(b-1)$ with $b-1\geq 0$, it follows that
$\dim\frac{S_{n-1}}{I_{t}(L_{n-1}^{s})}=n-1-a$
and thus
(3) $\operatorname{HS}(\frac{S}{I_{t}(L_{n-1})^{s}},z)=\frac{1}{1-z}\operatorname{HS}(\frac{S_{n-1}}{I_{t}(L_{n-1})^{s}},z)=\frac{1}{(1-z)^{n-a}}\mathbf{Q}(\frac{S_{n-1}}{I_{t}(L_{n-1}^{s})},z).$
Here, we use the assumption that $b\geq 1$. Note that

### v2

@@ 2. Associated primes of $I_{t}(L_{n})^{s}$
Let $G=\{a_{1},a_{2},\ldots,a_{2p}\}$. Then $n\geq a_{2p}=(t+1)p\geq n+t-1$ and $a_{2p-1}=t(p-1)+p<n-t+1$, and so it is routine to verify that $G\in\mathbf{C}_{n,t}$. This proves $D\geq 2p$. To prove $D\leq 2p$, we assume on the contrary that there exists $F=\{i_{1},\ldots,i_{r}\}\in\mathbf{C}_{n,t}$ with $1\leq i_{1}<\cdots<i_{r}\leq n$ such that $r>2p$. Then $i_{r-1}\geq i_{2p}\geq a_{2p}\geq n-t+1$. This is contradicted to condition (5). Hence, $D=2p$.
Case 2: If $n=(t+1)p+t$, then $n-t+1=(t+1)p+1=a_{2p+1},$ and one easily checks that $\{a_{1},\ldots,a_{2p+1}\}$ belongs to $\mathbf{C}_{n,t}$. It follows that $D=2p+1$, by using the same argument as in case 1.
The formula for the projective dimension of $S/I_{t}(L_{n})$ was first established in [15, Theorem 4.1], and subsequently proved through an alternative method in [5, Corollary 5.1]. In what follows, utilizing the theory of Alexander duality, we present a novel and concise demonstration. Recall a well-known result stating that $S/I$ is Cohen-Macaulay if and only if the Alexander dual $I^{\vee}$ of $I$ has a linear resolution, (see for example [12, Theorem 8.1.9]). Furthermore, according to [12, Theorem 8.2.20], $S/I$ is sequentially Cohen-Macaulay precisely when $I^{\vee}$ is componentwise linear. Finally, for any square-free monomial ideal $I$ of $S$, we have $\operatorname{pd}(S/I)=\operatorname{reg}(I^{\vee})$, as stated in [12, Proposition 8.1.10]. Here, $\operatorname{pd}$ refers to the projective dimension and $\operatorname{reg}$ refers to the (Castelnuovo-Mumford) regularity. For more about Alexander duality, we refer to Chapter 8 of the book [12]. .
@@
In [11, Corollary 5.4], it was shown that if $\Delta$ is a simplicial forest then $S/I(\Delta)$ is sequentially Cohen-Macaulay. Applying this result and in view of Lemma 2.2, we can conclude that $S/I_{t}(L_{n})$ is sequentially Cohen-Macaulay.
By [12, Theorem 8.2.20], we know $I_{t}(L_{n})^{\vee}$ is componentwise linear. Therefore, using [12, Corollary 8.2.14], we have $\operatorname{reg}(I_{t}(L_{n})^{\vee})=\operatorname{Deg}(I_{t}(L_{n})^{\vee})$. Furthermore, according to [12, Proposition 8.1.10], we have $\operatorname{pd}(S/I_{t}(L_{n}))=\operatorname{Deg}(I_{t}(L_{n})^{\vee})$. The result now follows from Corollary 2.9.
@@ 3. Multiplicity
In this section, we present the main result of this paper, which provides an explicit formula for determining the multiplicity of powers of $I_{t}(L_{n}).$ We will maintain all notation introduced in the previous sections. To start, we introduce a general result regarding the multiplicity of square-free monomial ideals.
@@
This follows directly from [12, Corollary 6.2.3].
@@
Assume that $n=\xi t+\theta$ with $\xi\geq 1$ and $0\leq\theta<t$. Then
$\operatorname{height}(I_{t}(L_{n}))=\xi.$
@@
We next describe the minimal prime ideals of $I_{t}(L_{n})$ of height $\xi$.
@@
Assume $n=\xi t+\theta$ with $\xi\geq 1$ and $0\leq\theta<t$. Let $\operatorname{\mathfrak{p}}=(x_{i_{1}},\ldots,x_{i_{\xi}})$ be a prime ideal of height $\xi$ with $1\leq i_{1}<\cdots<i_{\xi}\leq n$. Then $\operatorname{\mathfrak{p}}\in\mathrm{ass}(I_{t}(L_{n}))$ if and only if the following conditions are satisfied:
@@
Assume that the elements of $F=\{i_{1},\ldots,i_{\xi}\}$ meet conditions (1) and (2). We proceed to prove $F\in\mathbf{C}_{n,t}$, as defined in Notation 2.6. First, it is easy to see that $i_{1}\leq t$ and $i_{2}\geq(2-1)t+\theta+1\geq t+1$. Next, we observe that $i_{\xi-1}\leq(\xi-1)t=n-t-\theta<n-t+1$ and $i_{\xi}\geq(\xi-1)t+\theta+1=n-t+1$. Finally, for $j=1,\ldots,\xi-2$, we see that
$i_{j+2}-i_{j}\geq(j+1)t+\theta+1-jt=t+\theta+1\geq t+1.$
Thus, $F\in\mathbf{C}_{n,t}$ and so $\operatorname{\mathfrak{p}}\in\mathrm{ass}(I_{t}(L_{n}))$ by Proposition 2.7.
Conversely, since $\operatorname{\mathfrak{p}}\in\mathrm{ass}(I_{t}(L_{n}))$, we have $\{i_{1},\ldots,i_{\xi}\}$ is a minimal vertex cover of $\Delta_{n,t}$. Given that the facets $F_{1},F_{t+1},\ldots,F_{(\xi-1)t+1}$ and $F_{\theta+1},F_{t+\theta+1},\ldots,F_{(\xi-1)t+\theta+1}$ of $\Delta_{n,t}$ are pairwise disjoint within each respective set, it follows that for $j=1,\ldots,\xi$, one has $i_{j}$ belongs to both $F_{(j-1)t+1}$ and $F_{(j-1)t+\theta+1}$, and thus $(j-1)t+\theta+1\leq i_{j}\leq jt$, satisfying (1). As to condition (2), it follows directly from Proposition 2.7.
To determine the multiplicity of $S/I_{t}(L_{n})$, in light of Lemma 3.1, we must count the number of the sequences $(i_{1},\ldots,i_{\xi})$ that meet conditions (1) and (2), as described in Lemma 3.3. This counting problem is purely combinatorial and will be solved in the next two results. Note that for a finite set $A$, $|A|$ denotes the cardinal number of $A$, and for a positive integer $m$, $[m]$ denotes the set $\{1,\ldots,m\}$.
@@
Given $\xi\geq 1$ and $t\geq 1$, let $T_{\xi,t}$ denote the set of integral sequences $i_{1}<\cdots<i_{\xi}$ such that
@@
$|T_{\xi,t}|=\begin{pmatrix}\xi+t-1\\ \xi\end{pmatrix}.$
@@
We proceed by induction on $t$. If $t=1$, the result follows easily. Assume now that $t\geq 2$. For each sequence $(i_{1},i_{2},\ldots,i_{\xi})\in T_{\xi,t}$, we define
$\tau(i_{1},\ldots,i_{\xi}):=\operatorname{min}\{j\in[\xi]\colon\;i_{j}=(j-1)t+1\}.$
By convention, we set $\tau(i_{1},\ldots,i_{\xi})=\xi+1$ if $i_{j}\neq(j-1)t+1$ for all $j=1,\ldots,\xi$. Thus, $\tau(i_{1},\ldots,i_{\xi})\in\{1,\ldots,\xi,\xi+1\}$. We make the following easy observation:
if $\tau(i_{1},\ldots,i_{\xi})=k<\xi$, then $i_{j}=(j-1)t+1$ for $j=k,\ldots,\xi$.
For $k\in[\xi+1]$, we define
$T_{\xi,t}^{k}=\{(i_{1},i_{2},\ldots,i_{\xi})\in T_{\xi,t}\colon\;\tau(i_{1},\ldots,i_{\xi})=k\}.$
One easily sees that $|T_{\xi,t}^{1}|=1$ and $|T_{\xi,t}^{2}|=|\{i_{1}:2\leq i_{1}\leq t\}|=t-1$.
In general, for all $k\in[\xi+1]$, we use $T_{k}$ to denote the set of all sequences $i_{1}<i_{2}<\cdots<i_{k-1}$ such that $(j-1)t+2\leq i_{j}\leq jt$ for $j=1,\ldots,k-1,$ and $i_{j+1}-i_{j}\leq t$ for $j=1,\ldots,k-2.$ It follows from the above observation that $|T_{\xi,t}^{k}|=|T_{k}|$ for $k=1,\ldots,\xi+1$.
@@
$(i_{1},\ldots,i_{k-1})\longmapsto(i_{1}^{\prime},\ldots,i_{k-1}^{\prime})$
establishes a bijection between $T_{k}$ and $T_{k-1,t-1}$. Hence, $|T_{k}|$ is equal to the number of $T_{k-1,t-1}$. By the inductive hypothesis, we have $|T_{k-1,t-1}|=\begin{pmatrix}k+t-3\\ k-1\end{pmatrix}.$ Hence, it follows that
$|T_{\xi,t}|$ $=\mathop{\sum}\limits_{k=1}^{\xi+1}|T^{k}_{\xi,t}|=\mathop{\sum}\limits_{k=1}^{\xi+1}\begin{pmatrix}k+t-3\\ k-1\end{pmatrix}$
$=\begin{pmatrix}t-2\\ 0\end{pmatrix}+\begin{pmatrix}t-1\\ 1\end{pmatrix}+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}\xi+t-3\\ \xi-1\end{pmatrix}+\begin{pmatrix}\xi+t-2\\ \xi\end{pmatrix}$
$=\begin{pmatrix}t-1\\ 0\end{pmatrix}+\begin{pmatrix}t-1\\ 1\end{pmatrix}+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}\xi+t-3\\ \xi-1\end{pmatrix}+\begin{pmatrix}\xi+t-2\\ \xi\end{pmatrix}$
$=\begin{pmatrix}t\\ 1\end{pmatrix}+\begin{pmatrix}t\\ 2\end{pmatrix}+\cdots+\begin{pmatrix}\xi+t-3\\ \xi-1\end{pmatrix}+\begin{pmatrix}\xi+t-2\\ \xi\end{pmatrix}$
$=\begin{pmatrix}\xi+t-2\\ \xi-1\end{pmatrix}+\begin{pmatrix}\xi+t-2\\ \xi\end{pmatrix}$
$=\begin{pmatrix}\xi+t-1\\ \xi\end{pmatrix}.$
@@
Given $n\geq t\geq 1$ and assume $n=\xi t+\theta$ with $10\leq\theta\leq t-$ 1. Let $X_{n,t}$ denote the set of integral sequences $i_{1}<\ldots<i_{a}$ such that
@@
$|X_{n,t}|=\begin{pmatrix}\xi+t-\theta-1\\ \xi\end{pmatrix}.$
@@
Set $i_{j}^{\prime}=i_{j}-\theta j$. Then, we have for $1\leq j\leq\xi$,
$(j-1)t+\theta+1\leq i_{j}\leq jt\Longleftrightarrow(j-1)(t-\theta)+1\leq i_{j}^{\prime}\leq j(t-\theta),$
and for $11\leq j\leq\xi-$ 1,
$i_{j+1}-i_{j}\leq t\Longleftrightarrow i_{j+1}^{\prime}-i_{j}^{\prime}\leq t-\theta.$
Thus, the map: $(i_{1},\ldots,i_{\xi})\longmapsto(i_{1}^{\prime},\ldots,i_{\xi}^{\prime})$ establishes a bijection between $X_{n,t}$ and $T_{\xi,t-\theta}$. This implies
$|X_{n,t}|=|T_{\xi,t-\theta}|=\begin{pmatrix}\xi+t-\theta-1\\ \xi\end{pmatrix},$
@@
Assume $n=\xi t+\theta$ with $\xi\geq 1$ and $0\leq\theta<t$. Then
$\operatorname{mult}(S/I_{t}(L_{n}))=\begin{pmatrix}\xi+t-\theta-1\\ \xi\end{pmatrix}.$
@@
This is a combination of Lemma 3.1, Lemma 3.3, and Corollary 3.5.
Based on Proposition 3.6, we can obtain our main result through an induction process. In this process, constructing suitable short exact sequences is the crucial step. In order to achieve this, we require the following lemmas.
@@
The equality that $B_{i}=(I_{t}(L_{n-i})^{s},x_{n-i+1})$ is easy. Let us show that $A_{i}:x_{n-i+1}=A_{i+1}$. Since
@@
Let $T$ be a finitely generated graded $S$-modules of dimension $d$. According to [8, Corollary 4.1.8], we can express the Hilbert series of $T$ as
$\operatorname{HS}(T,z)=\frac{Q(T,z)}{(1-z)^{d}}$
for some $Q(T,z)\in\mathbb{Q}[z,z^{-1}]$ with $Q(T,1)\neq 0$. Furthermore, [8, Proposition 4.1.9] asserts
$\operatorname{mult}(T)=Q(T,1).$
Based on these results, we can promptly derive the subsequent lemma.
Lemma 3.9.
Let $0\longrightarrow K\longrightarrow M\longrightarrow N\longrightarrow 0$ be a short exact sequence of finitely generated graded $S$-modules. Then $\dim M=\operatorname{max}\{\dim K,\dim N\}$. Moreover,
  1. (1)
    If $\dim M=\dim K=\dim N$, then $\operatorname{mult}(M)=\operatorname{mult}(N)+\operatorname{mult}(K)$;
  2. (2)
    If $\dim M=\dim K>\dim N$, then $\operatorname{mult}(M)=\operatorname{mult}(K)$;
  3. (3)
    If $\dim M=\dim N>\dim K$, then $\operatorname{mult}(M)=\operatorname{mult}(N)$.
We also need the following fact: given that $I$ is a monomial ideal of $S$ and $y$ is a new variable, then
$\operatorname{mult}(S/I)=\operatorname{mult}(S[y]/IS[y]).$
@@
Assume $n=\xi t+\theta$, with $\xi\geq 1$ and $0\leq\theta<t$. Then, for all $s\geq 1$, we have
$\operatorname{mult}(S/I_{t}(L_{n})^{s})=\begin{pmatrix}s+\xi-1\\ s-1\end{pmatrix}\begin{pmatrix}\xi+t-\theta-1\\ \xi\end{pmatrix}.$
@@
It follows that $Q(S/I_{t}(L_{t})^{s},z)=\sum_{i=0}^{st-1}z^{i}$ and so $\operatorname{mult}(S/I_{t}(L_{t})^{s})=Q(S/I_{t}(L_{t})^{s},1)=st.$ This proves the result holds when $n=t$. Now, assume $s>1$ and $n>t$. To begin, we write $n=\xi t+\theta$, where $\xi\geq 1$ and $0\leq\theta<t$. There are two cases to consider.
Case 1: $\theta>0$. Consider the following three short exact sequences of graded $S$-modules:
(1) $0\longrightarrow\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}}[-t]\longrightarrow\frac{S}{I_{t}(L_{n})^{s}}\longrightarrow\frac{S}{(I_{t}(L_{n})^{s},u_{n-t+1})}\longrightarrow 0,$
@@
We proceed to show that all nonzero graded modules appearing in the above three short exact sequences share the same dimension, which is $(n-\xi)$. Firstly, by Lemma 3.7, we have $I_{t}(L_{n})^{s}:u_{n-t+1}=I_{t}(L_{n})^{s-1}$, it follows from Lemma 3.2 that
$\dim\frac{S}{I_{t}(L_{n})^{s}:u_{n-t+1}}=\dim\frac{S}{I_{t}(L_{n})^{s}}=n-\xi.$
Note that $(I_{t}(L_{n})^{s},u_{n-t+1})=(I_{t}(L_{n-1})^{s},u_{n-t+1})$ and $\sqrt{(I_{t}(L_{n-1})^{s},u_{n-t+1})}=I_{t}(L_{n})$, we obtain
$\dim\frac{S}{(I_{t}(L_{n})^{s},u_{n-t+1})}=\dim\frac{S}{(I_{t}(L_{n-1})^{s},u_{n-t+1})}=n-\xi.$
Since